OpenStudy (anonymous):

is an=4n+16 equal to any of these other options are an=16+(n-1)(4) an=16+4n an=20+(n)(4) an=20+(n-1)(4)

jimthompson5910 (jim_thompson5910):

Notice how an=20+(n-1)(4) turns into an=20+4n-4

jimthompson5910 (jim_thompson5910):

what does that simplify to?

OpenStudy (anonymous):

16+4n... but doesnt that mean it could be the first or last option @jim_thompson5910

jimthompson5910 (jim_thompson5910):

You mean the second and last?

OpenStudy (anonymous):

yes sorry @jim_thompson5910

jimthompson5910 (jim_thompson5910):

You're right, the second and last options are the same. So perhaps you can choose more than one answer?

jimthompson5910 (jim_thompson5910):

If you can, the answers are B and D

OpenStudy (anonymous):

well it says which ecxplicit formla corresponds with the simple so wouldnt you say D? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Yes, D is the basic form you start with

jimthompson5910 (jim_thompson5910):

Then with a bit of simplifying, it turns to B So if you can only choose one, I'd go with D